OpenStudy (anonymous):

The area of a triangle is 45.5 square meters and the base is 10 meters. What is the height of this triangle? A. 227.5 m B. 4.55 m C. 9.1 m D. 455 m

OpenStudy (gudden):

area of triangle = 1/2 (base ) (height) 45.5 = 1/2 (10)(height) on multiplying the equuation with 2 we get 91 = 10 ( height) so the height = 9.1 m

OpenStudy (a1234):

The formula for the area of a triangle is A = 1/2b*h. 45.5 = 1/2(10)*h 45.5 = 5h 9.1 = h

OpenStudy (anonymous):

lol ty guys

OpenStudy (gudden):

hah welcom:D

OpenStudy (a1234):

You're welcome.